# Multi-Type Point Cloud Autoencoder: A Complete Equivariant Embedding for Molecule Conformation and Pose

### v1

@@ Acknowledgements
The authors would like to thank Mihail Bogojeski for valuable discussions. JR acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG) through the Heisenberg Programme project 428315600. MK, JR, and MET acknowledge funding from grants from the National Science Foundation, DMR-2118890, and MET from CHE-1955381. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
Appendix A Supplemental material
A.1 Details of Model Architecture
@@
The EMLP layer acts on scalar inputs $\bm{s}$ of dimension $[n,k]$ and vector inputs $\bm{\vec{v}}$ of dimension $[n,3,l]$. The scalar track is a standard neural network layer with linear transformation, normalization, activation, dropout, and residual connection, with optional dimension adjustment between layers,
$\bm{s^{t+1}}=\bm{W}_{2}^{t}\left(\bm{s}^{t}+\mathrm{D}(\sigma(\mathrm{N}(\bm{W}_{1}^{t}(\bm{s}^{t}\lVert\tilde{\bm{v}}^{t})+\bm{b}^{t})))\right),$ (6)
and scalarized representation of the vector track $\bm{\tilde{v}}=\mathcal{S}(\bm{\vec{v}})$ concatenated before the linear operation. $\bm{\tilde{v}}$ is computed following panel (c) of Figure A1, and encodes pose-invariant information from $\vec{v}$, including the vector norms and relative internal angles. Following again the insight of [24], we learn a set of internal equivariant axes (in practice, three axes) within the normalized vector directions via $\bm{W}\bm{\bar{\vec{v}}}$, with $\bm{\bar{\vec{v}}}=\frac{\bm{\vec{v}}}{\lVert\bm{\vec{v}}\rVert}$ and compute their overlaps to $\bar{\vec{v}}$. This comprises a transform from the original Cartesian basis to a learned internal basis which rotates equivariantly with the molecule. We then concatenate this to the vector norms and pass them through a linear layer, treating them as regular invariant scalars. The obvious alternative to this ‘scalarization’ scheme is to simply take the euclidean norm of $\bm{\vec{v}}$ and use it as the ‘scalar representation’ of $\bm{\vec{v}}$. While reasonable in principle, this leaves invariant information about the relative directions of vectors on the table, and so $\bm{\tilde{v}}$ should provide a richer scalar picture than $\bm{\bar{v}}$.
The vector side of the forward pass is analogous to the scalar, but with adjustments to retain equivariance.
$\begin{split}\bm{s_{g}}&=\sigma(\mathrm{N}(\bm{W}^{t}\bm{s}^{t+1}+\bm{b}_{s}^{t})),\\ \bm{\vec{v}}_{g}&=\vec{\sigma}(\vec{N}(\bm{W}^{t}\vec{v}^{t}))),\\ \bm{\vec{v}}^{t+1}&=\bm{W}^{t}\left(\vec{\bm{v}}^{t}+\bm{s_{g}}\odot\bm{\vec{v}}_{g}\right).\end{split}$ (7)
@@
Graph convolution again comprises a scalar and vector track, with the scalar track using the ‘TransformerConv’ module [60], an efficient and expressive graph convolution which combines neighboring nodes according to a coefficient $\alpha_{ij}$ learned via the self-attention mechanism,
$\alpha_{ij}=\mathrm{Softmax}\left(\frac{\left(\bm{W}_{1}\bm{S}_{i}\right)^{T}\left(\bm{W}_{2}\bm{S}_{j}+\bm{W}_{3}\bm{E}_{ij}\right)}{\sqrt{d}}\right),$ (9)
with $\bm{S}_{i}=\bm{W}\bm{s}^{t}$ and $\bm{E}_{ij}=\bm{W}\bm{e}_{ij}$ as node and edge embeddings for message passing respectively. Scalar messages are passed according to
$\bm{S}_{i}^{t+1}=\beta_{i}\bm{W}_{1}\bm{S}_{i}^{t}+\left(1-\beta_{i}\right)\sum_{j\in r_{c}}\alpha_{ij}\left(\bm{W}_{2}\bm{S}_{j}^{t}+\bm{W}_{3}\bm{E}_{ij}\right),$ (10)
with $j\in r_{c}$ the nodes $j$ within the cutoff range $r_{c}$ of node $i$. Vector messages are passed according to an equivariant attention function with shared attention weights from the scalar message passing, where we gate against the edge contributions rather than adding to maintain equivariance,
$\vec{\bm{V}}_{i}^{t+1}=\bm{W}_{1}\vec{\bm{V}}_{i}^{t}+\sum_{j\in r_{c}}\alpha_{ij}\left(\bm{W}_{2}\vec{\bm{V}}_{j}^{t}\cdot\bm{W}_{3}\bm{E}_{ij}\right).$ (11)
After graph convolution, we aggregate nodewise information via a softmax weighting over the channel dimension,
$\vec{\bm{g}}_{k}=\sum_{i}^{n_{i}}\left(\frac{e^{\bar{\bm{v}}_{ik}}}{\sum_{j}^{n_{i}}e^{\bar{\bm{v}}_{jk}}}\right)\vec{\bm{v}}_{ik}.$ (12)
@@
A.2 Model hyperparamters and optimization
@@
A single training run of one of our autoencoder models on QM9 took 1-2 days on an NVIDIA V100 GPU. A QM9 property regression run took 1-3 hours on the same hardware. Over the course of the project, several hundred partial and complete runs were undertaken for the development of the new architecture and loss model, and to identify hyperparameter regimes with fast convergence go high quality loss minima. We found by experimentation that different hyperparameters significantly impacted the generalization and training speed of the autoencoder model. Here, we give the best parameters found to-date, which were used for the results shown in the main text.
@@
Encoder: GeLU activation, atom type embedding dimension of 32, graph convolution cutoff of 14Å, 5% dropout probability in nodewise fully-connected layers, node embedding dimension, graph readout dimension and bottleneck dimension all 512, 4 fully-connected layers per EMLP block, message dimension of 128, 16 attention heads, graph-wise layer normalization, 1 convolutional layer, and 32 Gaussian radial basis functions.
Decoder: GeLU activation, 5% dropout probability, hidden dimension of 512, 512 nodes in the output swarm, and 1 fully-connected layer.
@@
Autoencoder Optimization: AdamW [61] optimizer with weight decay of 0.05, initial learning rate of 5.0E-5, ramping to 1.0E-3, then down to 1.0E-6. Batch size starting at 50 and ramping to 2522 (maxes out our GPU memory). The Gaussian width is initialized at 1.05, and reduced by 1% when the reconstruction test loss reaches 0.01, until the mean sample self-overlap reaches 5.0E-4.
@@
Embedding Regressor: GeLU activation, 10% dropout probability, hidden dimension of 128, and 4 fully-connected layers.

### v3

@@ Acknowledgements
The authors would like to thank Mihail Bogojeski for valuable discussions. JR acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG) through the Heisenberg Programme project 428315600. MK, JR, and MET acknowledge funding from grants from the National Science Foundation, DMR-2118890, and MET from CHE-1955381. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The Flatiron Institute is a division of the Simons Foundation.
Appendix A Details of Model Architecture
@@
The EMLP layer acts on scalar inputs $\bm{s}$ of dimension $[n,k]$, for $n$ the batch index and $k$ the feature depth for scalars, and vector inputs $\bm{\vec{v}}$ of dimension $[n,3,l]$, for $l$ the feature depth for vectors. The scalar track is a standard neural network layer acting on with linear transformation according to learned weight matrix $W^{t}$ and bias $b^{t}$, normalization $\mathrm{N}$, activation $\sigma$, dropout $\mathrm{D}$, and residual connection (addition by $\bm{s}^{t}$), with optional dimension adjustment between layers,
$\bm{s^{t+1}}=\bm{s}^{t}+\mathrm{D}(\sigma(\mathrm{N}(\bm{W}^{t}(\bm{s}^{t}+\tilde{\bm{v}}^{t})+\bm{b}^{t}))),$ (10)
and scalarized representation of the vector track $\bm{\tilde{v}}=\mathcal{S}(\bm{\vec{v}})$ added before the linear operation.
$\bm{\tilde{v}}$ is computed following panel (c) of Figure A1, and encodes pose-invariant information from $\vec{v}$, including the vector norms and relative internal angles. Following again the insight of [23], we learn a set of internal axes (in practice, three) within the normalized vector directions via $\bm{W}\bm{\bar{\vec{v}}}$, with $\bm{\bar{\vec{v}}}=\frac{\bm{\vec{v}}}{\lVert\bm{\vec{v}}\rVert}$ and compute their overlaps to $\bar{\vec{v}}$. This comprises a transform from the original Cartesian basis to a learned internal basis which rotates equivariantly with the molecule. We then concatenate this to the vector norms and pass them through a linear layer, treating them as regular invariant scalars. The obvious alternative to this ‘scalarization’ scheme is to simply take the euclidean norm of $\bm{\vec{v}}$ and use it as the ‘scalar representation’ of $\bm{\vec{v}}$. $\bm{\bar{v}}$ lacks invariant information about the relative directions of vectors on the table; $\bm{\tilde{v}}$ should provide a richer picture.
The vector side of the forward pass is analogous to the scalar, but with adjustments to retain equivariance,
$\begin{split}\bm{s_{g}}&=\sigma(\mathrm{N}(\bm{W}^{t}\bm{s}^{t+1}+\bm{b}_{s}^{t})),\\ \bm{\vec{v}}_{g}&=\vec{\sigma}(\vec{N}(\bm{W}^{t}\vec{v}^{t}))),\\ \bm{\vec{v}}^{t+1}&=\vec{\bm{v}}^{t}+\bm{s_{g}}\odot\bm{\vec{v}}_{g}.\end{split}$ (11)
@@
Aggregations during graph convolution and global aggregation are both handled by an augmented softmax operation. We add to the work of [71], wherein softmax aggregation was shown to smoothly interpolate between the standard ‘mean’ and ‘max’ aggregation operations via a channel-dependent learned temperature, $T_{k}=1/\beta_{k}$. We boost the flexibility of this approach with the addition of a second learned parameter, $b_{k}$, which adds ‘sum’ to the set of aggregation operations smoothly interpolated by the aggregator, which we call ‘augmented softmax’,
$\bm{g}_{k}=\sum_{i}^{n_{i}}\left(\frac{e^{\beta_{k}x_{ik}}}{\sum_{j}^{n_{i}}e^{\beta_{k}x_{jk}}}+b_{k}\right)x_{ik},$ (13)
where sum is taken over the number of nodes to be aggregated, $n_{i}$, $x_{ik}$ is the k-th channel feature for node $i$, and $\bm{g}_{k}$ the aggregated value for channel $k$.
We use this function for the aggregation of both scalars and vectors, with the modification that when aggregating vectors, the norm is substituted for the scalar in the softmax operation
$\vec{\bm{g}}_{k}=\sum_{i}^{n_{i}}\left(\frac{e^{\beta_{k}\bar{\bm{v}}_{ik}}}{\sum_{j}^{n_{i}}e^{\beta_{k}\bar{\bm{v}}_{jk}}}+b_{k}\right)\vec{\bm{v}}_{ik},$ (14)
now with $\vec{\bm{v}}_{ik}$ the vector-valued k-th channel feature for node $i$ and $\bar{\bm{v}}_{i}k$ its Euclidean norm.
Graph convolution again comprises a scalar and vector track, with scalar messages passed according to
$\bm{s}_{i}^{t+1}=\bm{s}_{i}^{t}+\bm{W}_{1}Agg_{j}(\sigma(N(\bm{W}_{2}\bm{s}_{i}^{t}|\bm{W}_{3}\bm{s}_{j}^{t}|\bm{W}_{4}\bm{e}_{ij}))),$ (15)
with $\bm{W}_{1}$-$\bm{W}_{4}$ learnable weight matrices, $\bm{e}_{ij}$ the feature vector for the $j\to i$ edge, $j\in r_{c}$ the nodes $j$ within the cutoff range $r_{c}$ of node $i$, and $Agg_{j}$ the augmented softmax aggregation over such nodes. Vector messages are passed analogously, except we gate against the edge contributions rather than adding to maintain equivariance,
$\vec{\bm{v}}_{i}^{t+1}=\vec{\bm{v}}_{i}^{t}+\bm{W}_{1}Agg^{v}_{j}(N(\bm{W}_{2}\vec{\bm{v}}_{i}^{t}+\bm{W}_{3}\vec{\bm{v}}_{j}^{t})\cdot\bm{W}_{4}\bm{e}_{ij}),$ (16)
with $Agg_{j}^{v}$ the vector augmented softmax aggregation.
@@
Appendix B Mean Deviation Calculation
For molecule reconstruction, the mean deviation between input and outputs is calculated via a scaffolding and averaging procedure. For each atom in the input point cloud (molecule), GM output components within 0.5 Å are collected, and a synthetic atom is generated by the weighted sum of their positions. If the total probability mass of the synthetic atom is above a certain cutoff, in this study, 0.25, we count the input atom as ‘matched’, and compute the distance between the two.
The average over all atoms in a dataset is done simply as $\Delta_{atomwise}=\frac{1}{n}\sum_{i}^{n_{i}}\delta_{i}$, with $\delta_{i}$ the computed deviation for atom $i$ and $n$ as the number of atoms in the dataset. The molecule-wise average is done instead by first averaging over the number of atoms in each molecule, and then the number of molecules in the dataset, as $\Delta_{molwise}=\frac{1}{n_{m}}\sum_{m}^{n_{m}}\frac{1}{n_{i}}\sum_{i}^{n_{i}}\delta_{i}$, for $n_{m}$ the number of molecules in the dataset, and $n_{i}$ the number of atoms in molecule $m$.
Appendix C Regression heads
Scalar regression and proxy discriminator regression heads are straightforward EMLP models with scalar outputs. The crystal embedding model is a simple scalar MLP.
For the regression of the QM9s vector properties, we use the vector output of the EMLP model.
For regression of rank 2 tensors (polarizability and quadrupole moment), we adopt the following ansatz: first split the scalar and vector output channels into two equally sized groups, $\bm{s}=\bm{s}_{1}\oplus\bm{s}_{2}$, $\vec{\bm{v}}=\vec{\bm{v}}_{1}\oplus\vec{\bm{v}}_{2}$, then construct a symmetric tensor via isotropic and anisotropic parts,
$\bm{T}^{(2)}=\sum_{i}^{n_{out}}\bm{s}_{1,i}\cdot\bm{I}^{3}+w_{i}\frac{1}{2}(\vec{\bm{v}}_{1,i}\otimes\vec{\bm{v}}_{2,i}+\vec{\bm{v}}_{2,i}\otimes\vec{\bm{v}}_{1,i}),$ (17)
where $w_{i}$ is a scalar weight given by a softmax on $s_{2}$, $\bm{I}^{3}$ is the identity matrix, and $n_{out}$ is the number of components in the output.
Rank 3 tensors are constructed analogously, but with the ranks increased by one, and the the output channels split into four batches,
$\bm{T}^{(3)}=\sum_{i}^{n_{out}}\frac{1}{3}(\bm{I}^{3}\bm{v}_{4,i}+(\bm{I}^{3}\bm{v}_{4,i})^{(13)}+(\bm{I}^{3}\bm{v}_{4,i})^{(12)})+(\vec{\bm{v}}_{1,i}\otimes\vec{\bm{v}}_{2,i}\otimes\vec{\bm{v}}_{3,i}+...+\vec{\bm{v}}_{3,i}\otimes\vec{\bm{v}}_{2,i}\otimes\vec{\bm{v}}_{1,i}),$ (18)
with the first sum over the outer products of identity matrices and input vectors with permuted indices, and the sum in the second brackets taken over the six permutations of the indices. The first term embeds vectors as 3-tensors, replacing the scaled identity matrix, and the second term constructs the full symmetric tensor.
Appendix D Model hyperparamters and optimization
@@
A single training run of one of our autoencoder models on QM9 took 2 days on an NVIDIA RTX8000 GPU. A QM9 property regression run took 1-3 hours on the same hardware, and a crystal energy regression takes 4-8 hours. Over the course of the project, several hundred partial and complete runs were undertaken for the development of the new architecture and loss model, and to identify hyperparameter regimes with fast convergence go high quality loss minima. We found by experimentation that different hyperparameters significantly impacted the generalization and training speed of the autoencoder model. Here, we give the best parameters found to-date, which were used for the results shown in the main text.
@@
Encoder: GeLU activation, atom type embedding dimension of 32, graph convolution cutoff of 3Å, no dropout, node embedding dimension of 512, bottleneck dimension of 64, 2 fully-connected layers per EMLP block, message dimension of 64, graph-wise layer normalization on scalars but not vectors, 2 convolutional layers, and 50 Gaussian radial basis functions.
Decoder: GeLU activation, no dropout, 64 nodes per graph, node dimension of 128, 4 convolutional layers, and graph-wise layer normalization on scalars but not vectors.
@@
Autoencoder Optimization: AdamW [72] optimizer with weight decay of 0.005, initial learning rate of 5.0E-5, ramping to 1.0E-3, then down to 1.0E-6. Batch size starting at 10 and ramping until GPU memory is maxed out at 167. The Gaussian width is initialized at 2, and reduced by 1% when the reconstruction test loss reaches 0.15, until the mean sample self-overlap reaches 1.0E-3.
@@
Embedding Regressor: GeLU activation, no dropout or normalization, hidden dimension of 256, and 4 fully-connected layers.
@@
Crystal Energy Model: GeLU activation, hidden dimension of 512, and 32 fully-connected layers. Layer normalization and 50% dropout probability.
Energy Model Optimization: AdamW optimizer with weight decay of 0.1, initial learning rate of 1.0E-4, ramping to 5.0E-4, then down to 5.0E-6. Batch size starting at 50 and ramping to 10000.
Appendix E Molecule property prediction
For reference, we provide in Figures 914 a more detailed view of the regression results for the QM9/QM9s molecular properties.
Figure 9: Regression results for rotational spectroscopic constants.
Figure 10: Regression results for electronic properties.
Figure 11: Regression results for molecular orbital energies.
Figure 12: Regression results for zero point vibrational and internal energies.
Figure 13: Regression results for enthalpy, free energy, and heat capacity.
Figure 14: Regression results for QM9s vector and tensor properties.
Appendix F Crystal Parameterization
Figure 15: Molecular crystal parameterization, construction, and scoring. The molecule comprising the asymmetric unit is taken in a given pose, reoriented by an axis-angle rotation vector $\mathbf{\vec{\theta}}$ and placed at position $\mathbf{\vec{T}}$ in the unit cell defined by $\mathbf{\hat{H}}$. The symmetry operations $\mathbf{\hat{S}}$ are applied depending on the relevant space group, and a cluster of molecules within the intermolecular interaction radius of the asymmetric unit is added. The intermolecular energy is computed on edges between the asymmetric unit, highlighted in green in panel (e), within a range of 6 Å. We show here a molecule in space group 14 ($P2_{1}/c$), for illustration. Actual training runs were all done in space group 1 ($P$1), where the asymmetric unit is identical to the unit cell.
Molecular crystals with $Z^{\prime}=1$ (one molecule per asymmetric unit) can be completely characterized by the molecule point cloud, the space group, and the 12 crystal dimensions: cell vector lengths ($\bar{a}$,$\bar{b}$,$\bar{c}$), internal angles ($\alpha$, $\beta$, $\gamma$), the position of the molecule centroid in fractional coordinates $\vec{T}=(u,v,w)$, and the molecule orientation defined against a standardized orientation, defined by a rotation vector $\vec{(}\theta)=(x,y,z)$. The chosen standard is to align the molecule principal eigenvectors to the cartesian axes.
Lennard-Jones potentials for each molecular crystal are computed over intermolecular edges as
$E_{LJ}=4\sum_{ij}\left(\frac{\sigma_{ij}}{r_{ij}^{12}}-\frac{\sigma_{ij}}{r_{ij}^{6}}\right),$ (19)
for $\sigma_{ij}$ the sum of atoms $i$ and $j$’s van der Waals radii.